# FLoRA: Low-Rank Core Space for N-dimension

### v1

@@ Abstract
Adapting pre-trained foundation models for various downstream tasks has been prevalent in artificial intelligence. Due to the vast number of tasks and high costs, adjusting all parameters becomes unfeasible. To mitigate this, several fine-tuning techniques have been developed to update the pre-trained model weights in a more resource-efficient manner, such as through low-rank adjustments. Yet, almost all of these methods focus on linear weights, neglecting the intricacies of parameter spaces in higher dimensions like 4D. Alternatively, some methods can be adapted for high-dimensional parameter space by compressing changes in the original space into two dimensions and then employing low-rank matrix decomposition. However, these approaches destructs the structural integrity of the involved high-dimensional spaces. To tackle the diversity of dimensional spaces across different foundation models and provide a more precise representation of the changes within these spaces, this paper introduces a generalized parameter-efficient fine-tuning framework, FLoRA, designed for various dimensional parameter space. Specifically, utilizing Tucker decomposition, FLoRA asserts that changes in each dimensional parameter space are based on a low-rank core space which maintains the consistent topological structure with the original space. It then models the changes through this core space alongside corresponding weights to reconstruct alterations in the original space. FLoRA effectively preserves the structural integrity of the change of original N-dimensional parameter space, meanwhile decomposes it via low-rank tensor decomposition. Extensive experiments on computer vision, natural language processing and multi-modal tasks validate FLoRA’s effectiveness. Codes are available at https://github.com/SJTU-DeepVisionLab/FLoRA.
@@ 1 Introduction
Figure 1: Difference between LoRA and FLoRA. LoRA employs low-rank matrix decomposition for each dimensional parameter space. However, for parameter space of convolution layer, the reshaping operation required by LoRA causes adjacent elements within the kernel to be separated in the reshaped matrix, disrupting the spatial locality inherent in the original convolutional space. Conversely, FLoRA asserts that the alternations of each dimensional parameter space has a low-rank core space with the consistent topological structure. This framework enables FLoRA to effectively preserve the structural integrity of the original parameter space, such as maintaining the spatial locality in convolutional operations.
To tackle this challenge, recent works Chen et al. (2024); Guo et al. (2020); He et al. (2021a); Hu et al. (2021) have focused on the concept of parameter-efficient fine-tuning (PEFT), aiming to minimize the number of adjustable parameters required while achieving optimal task performance. These works specifically explore methods to model the incremental update of pre-trained weights in a manner that is economical in terms of parameters, without necessitating changes to the model’s architecture Zaken et al. (2021); Guo et al. (2020); Hu et al. (2021). Among these works, LoRA Hu et al. (2021) emerges as a pioneering effort, which proposes to adopt an additional term with a low-rank structure to the original weight. Specifically, the original weight matrix $\mathbf{W}_{0}\in\mathbb{R}^{d_{1}\times d_{2}}$ remains frozen, while a learnable low-rank $\Delta\mathbf{W}$ is added to $\mathbf{W}_{0}$, with the form as
@@
where $\mathbf{A}\in\mathbb{R}^{d_{2}\times r}$ and $\mathbf{B}\in\mathbb{R}^{d_{1}\times r}$ with $r\ll\{d_{1},d_{2}\}$. Given that $r$ is significantly smaller than the dimension of $\mathbf{W}_{0}$, LoRA necessitates the updating of only a limited number of trainable parameters, while achieving comparable performances. Subsequent to LoRA, recent studies He et al. (2021a); Bershatsky et al. (2024); Zhang et al. (2022) endeavor to explore more efficient low-rank matrix decomposition methods concerning the matrix $\Delta\mathbf{W}$.
However, we have observed an intriguing phenomenon: a significant portion of existing works is narrowly focused on two-dimensional parameter spaces (i.e., linear layers), neglecting the existence of other dimensional spaces such as 2D convolution (4-dimension) layers, etc. But in practice, not all layers are linear across various models for downstream tasks. For example, ConvNeXt Liu et al. (2022) and Stable Diffusion Carreira and Zisserman (2017) are two models predominantly utilizing 2D convolutional layers. Alternatively, some approaches can be adapted for high-dimensional parameter spaces by directly adopting low-rank matrix decomposition. They transform alternations in high-dimensional spaces into two dimensions, while neglecting the structural complexities of the original spaces. For instance, as detailed in Sec. 3.1 and Fig. 1, LoRA Hu et al. (2021) compresses the changes in convolutions, which is a four-dimensional parameter space, into two dimensions. It subsequently applies low-rank matrix decomposition to the two-dimension space, , intending to represent the alterations of the original four-dimensional parameter space. Yet, as discussed in Sec. 3.2, this transformation fails to capture the inherent complexity and spatial locality specific to convolution operations. The result is a reshaped two-dimensional structure that compromises the integrity of the original parameter space, leading to a representation that does not fully encapsulate the convolutional alternation parameter space.
To this end, in this paper, we propose a low-rank tensor decomposition based method, FLoRA, represented as Fundamental LOw-Rank Adaptation. Positioned as a superior alternative to LoRAs, FLoRA, based on Tucker decomposition Tucker (1966), meets the ensuing three properties:
  • It can identify a appropriate low-rank representation for the changes in various dimensional parameter spaces, without destructing the structural integrity of the original parameter spaces.
@@
  • When applied to linear weights, with the same parameter budget, it requires a similar training time and resources to that of LoRAs, yet yield superior performance.
@@
2 Related Work
Methods for Parameter-Efficient Fine-tuning (PEFT) have been conceived to mitigate the substantial computational costs associated with the fine-tuning of large-scale models. This economization is realized by honing a comparatively minute fraction of the overall parameters, selected strategically for adaptation to a variety of downstream tasks. Current PEFT techniques can be divided into three distinct categories Liu et al. (2024b); Ding et al. (2023): Adapter-based Houlsby et al. (2019); Chen et al. (2022); Luo et al. (2023); He et al. (2021a); Mahabadi et al. (2021); Karimi Mahabadi et al. (2021), prompt-based Lester et al. (2021); Razdaibiedina et al. (2023); Wang et al. (2023b); Shi and Lipani (2023); Fischer et al. (2024) and low-rank matrix decomposition-based Hu et al. (2021); Liu et al. (2024b); Hyeon-Woo et al. (2021); Qiu et al. (2023); Renduchintala et al. (2023); Kopiczko et al. (2023); YEH et al. (2023); Zhang et al. (2022). The first category of method integrates linear modules either sequentially or concurrently with the existing layer to enhance the performance, and the second class introduces additional soft tokens (prompts) to the initial input and concentrate exclusively on refining these trainable vectors. The last type, proposed by LoRA Hu et al. (2021), adopts low-rank matrix decomposition to model the change of the weight during fine-tuning, and are capable of merging with pre-trained weights.
However, as mentioned in the introduction, these methods only focuses on linear weight or destructing the structural integrity of high dimensional parameter spaces. To this end, in this paper we propose a novel PEFT method, aimed to address various dimensional parameter space.
@@ 3.1 Low Rank Adaptation
LoRA Hu et al. (2021) models the incremental update of a pre-trained weight matrix $\mathbf{W}_{0}\in\mathbb{R}^{d_{1}\times d_{2}}$ by the product of two small matrices $\mathbf{A}\in\mathbb{R}^{d_{2}\times r}$ and $\mathbf{B}\in\mathbb{R}^{d_{1}\times r}$, where $r\ll\{d_{1},d_{2}\}$. For $\mathbf{h}=\mathbf{W}_{0}\mathbf{x}$, the modified forward pass is
@@
where $\mathbf{A}_{c}\in\mathbb{R}^{kd_{out}\times r}$ and $\mathbf{B}_{c}\in\mathbb{R}^{kd_{in}\times r}$. Here, $Reshape(\mathbf{B}_{c}\mathbf{A}_{c}^{\mathsf{T}},\mathcal{W}_{c})$ involves altering the dimensions of $\mathbf{B}_{c}\mathbf{A}_{c}^{\mathsf{T}}$ to match those of $\mathcal{W}_{c}$. It is obvious that LoRA unfolds the original 4-dimensional parameter space $\Delta\mathcal{W}$ into a 2-dimensional space, subsequently leveraging the low-rank approximation of this 2-dimensional space to represent the original 4-dimensional construct. As shown later, this low-rank matrix decomposition based method will destruct the structural integrity of the convolution layer.
3.2 Why Matrix Decomposition Breaks the Structural Integrity of the Convolution?
A high-dimension tensor can be decomposed based on low-rank matrix decomposition following Eq. (3). However, during the reshaping process, elements that were adjacent within the convolutional kernel may become scattered across various positions in the matrix. More specifically, elements that were localized within the convolutional kernel may now span multiple rows or columns of the reshaped matrix. This shift poses significant challenges in learning spatial correlations among elements positioned disparately. Therefore, such a transformation disrupts the principle of locality inherent in the original convolutional operation, where each output element is determined by a small region of the input.
3.3 Tucker Decomposition
Tucker decomposition Tucker (1966) is one of the well-studied algebraic tensor decomposition. Formally, given a tensor $\mathcal{X}\in\mathbb{R}^{I_{1}\times I_{2}\times\cdots\times I_{N}}$, where $N$ is the order of the tensor (i.e., the number of dimensions or modes), Tucker decomposition represents $\mathcal{X}$ as a product of a core tensor $\mathcal{G}\in\mathbb{R}^{J_{1}\times J_{2}\times\cdots\times J_{N}}$ and a matrix along each mode $n$, $\mathbf{A}^{(n)}\in\mathbb{R}^{I_{n}\times J_{n}}$, where $J_{n}$ can be considered as the dimension of the core tensor along mode $n$. The decomposition can be compactly written as:
where ×$\times_{n}$ denotes the or mode-$n$ product between a tensor and a matrix. The core tensor $\mathcal{G}$ represents the interactions between different modes, and the matrices $\mathbf{A}^{(n)}$ are analogues to principal components within each respective mode. The selection of dimensions $J_{1},J_{2},\ldots,J_{N}$ allows for a balance between desired approximation quality and computational efficiency, tailored to the specific requirements of the task at hand. Tucker decomposition decomposes any dimensional space into a core tensor and a set of matrices related to the dimensions, without altering the original structure. Specifically, these matrices only transform the representation of different dimensions, which means that as long as the core tensor accurately captures the relationships among these dimensions, the property of original parameter space can be preserved through this decomposition.
@@ 4 Method
In this section, we first introduce the formulation of FLoRA in an N-dimensional parameter space. Specifically, for a pre-trained weight $\mathcal{W}_{0}\in\mathbb{R}^{I_{1}\times I_{2}\times\cdots\times I_{N}}$ with N-dimension, we model its update $\Delta\mathcal{W}$ as
@@
without loss of generality. Here $\mathcal{G}\in\mathbb{R}^{J_{1}\times J_{2}\times\cdots\times J_{N}}$ and $\mathbf{A}^{(n)}\in\mathbb{R}^{I_{n}\times J_{n}}$, where $J_{n}\ll I_{n}$. FLoRA considers the tensor $\mathcal{G}$ as a low-rank core space with the consistent topological structure to the original parameter space, with $\mathbf{A}^{(n)}$ representing the weights associated with each dimension. We then scale $\Delta\mathcal{W}$ by $s$ with $s$ being a constant. In the subsequent subsections, we will detail its manifestations in convolution and linear layers.
@@ 4.1 FLoRA for Convolution Layer
where $\mathcal{G}\in\mathbb{R}^{r_{1}\times r_{2}\times r_{3}\times r_{3}}$, $\mathbf{A}\in\mathbb{R}^{d_{in}\times r_{1}}$, $\mathbf{B}\in\mathbb{R}^{d_{out}\times r_{2}}$ and $\mathbf{C},\mathbf{D}\in\mathbb{R}^{k\times r_{3}}$. $r_{1},r_{2}$ are ranks which are significantly smaller than $\{d_{in},d_{out}\}$. $r_{3}$ is a rank smaller than the kernel size $k$ of a convolution layer. Given that $3\times 3$ convolution is a prevalent configuration in convolutional foundation models Woo et al. (2023); Rombach et al. (2022); Wang et al. (2023a), $r_{3}$ is consequently set to {1, 2}.
The core tensor $\mathcal{G}$ in FLoRA can be viewed as a compressed convolution parameter space. In essence, it serves as a core space for convolution. This means that in any convolution layer, there exists a convolution core, and what FLoRA aims to do is to determine this convolution core for each convolution space and configure the corresponding weights $\mathbf{A}$, $\mathbf{B}$, $\mathbf{C}$ and $\mathbf{D}$ for reconstructing that space. Different to low-rank matrix decomposition based methods, FLoRA dose not need to alter the structure of the convolution. Alternatively, by learning the convolution core, FLoRA effectively preserves the convolution’s property of spatial locality.
Furthermore, while preserving or potentially augmenting the representational power of the convolution process, FLoRA achieves a remarkable reduction in the number of trainable parameters in comparison to LoRAs. Assuming that the rank for both the input and output dimensions is uniform ($r_{1}=r_{2}=r$), the parameter requirement for FLoRA is calculated as $O(r_{3}^{2}r^{2}+r(d_{in}+d_{out})+2r_{3}k)$ parameters, while LoRAs necessitate training at least $O(rk(d_{in}+d_{out}))$ parameters. Given that typically $r,k\ll\{d_{in},d_{out}\}$, therefore, FLoRA exhibits better parameter efficacy than LoRA as the number of the kernel increases.
@@ 5.1 Models and Datasets
Specifically, for CV tasks, we employ FLoRA to fine-tune the ConvNeXt-V2-L Woo et al. (2023), evaluating it on MS COCO Lin et al. (2014) by using Mask R-CNN He et al. (2017) implemented in MMDetection Chen et al. (2019), and on remote sensing image datasets DOTA Xia et al. (2018) with Oriented R-CNN Xie et al. (2021) based on MMRotate Zhou et al. (2022), and on the ADE20K Zhou et al. (2017) dataset thanks to UperNet Xiao et al. (2018) integrated in MMSegmentation Contributors (2020). We also employ FLoRA to fine-tune large-scale vision foundation model, i.e. InternViT-6B Chen et al. (2023), on ADE20K datasets. Detailed hyper-parameter settings can be find in Appendix.
For NLP tasks, we evaluate the DeBERTaV3-base He et al. (2021b) with FLoRA on the General Language Understanding Evaluation (GLUE) Wang et al. (2018) benchmark, which includes two single-sentence classification, three similarity and paraphrase and four natural language inference datasets. More details on GLUE dataset can be found in the Appendix.
For multi-modal tasks, we employ FLoRA to fine-tune LLaVA-1.5-7B Liu et al. (2024a), which is composed of a language model, Vicuna-1.5-7B Peng et al. (2023) and a vision encoder, CLIP ViT-L/336px Radford et al. (2021), on visual instruction tuning tasks, which include seven vision-language benchmarks: VQAv2 Goyal et al. (2017), GQA Hudson and Manning (2019), VisWiz Gurari et al. (2018), SQA Lu et al. (2022), VQAT Singh et al. (2019), POPE Li et al. (2023), and MMBench Liu et al. (2023).
Moreover, we have also fine-tune Stable Diffusion Rombach et al. (2022) with FLoRA in Appendix.

### v2

# Maintaining Structural Integrity in Parameter Spaces for Parameter Efficient Fine-tuning

@@ Abstract
Adapting pre-trained foundation models for various downstream tasks has been prevalent in artificial intelligence. Due to the vast number of tasks and high costs, adjusting all parameters becomes unfeasible. To mitigate this, several fine-tuning techniques have been developed to update the pre-trained model weights in a more resource-efficient manner, such as through low-rank adjustments. Yet, almost all of these methods focus on linear weights, neglecting the intricacies of parameter spaces in higher dimensions like 4D. Alternatively, some methods can be adapted for high-dimensional parameter space by compressing changes in the original space into two dimensions and then employing low-rank matrix adaptations. However, these approaches destructs the structural integrity of the involved high-dimensional spaces. To tackle the diversity of dimensional spaces across different foundation models and provide a more precise representation of the changes within these spaces, this paper introduces a generalized parameter-efficient fine-tuning framework, designed for various dimensional parameter space. Specifically, our method asserts that changes in each dimensional parameter space are based on a low-rank core space which maintains the consistent topological structure with the original space. It then models the changes through this core space alongside corresponding weights to reconstruct alterations in the original space. It effectively preserves the structural integrity of the change of original N-dimensional parameter space, meanwhile models it via low-rank tensor adaptation. Extensive experiments on computer vision, natural language processing and multi-modal tasks validate the effectiveness of our method.
@@ 1 Introduction
Figure 1: Difference between LoRA and FLoRA. LoRA employs low-rank matrix adaptation for each dimensional parameter space. However, for parameter space of convolution layer, the reshaping operation required by LoRA causes adjacent elements within the kernel to be separated in the matrix, disrupting the spatial locality inherent in the original convolutional space. Conversely, FLoRA asserts that the alternations of each dimensional parameter space has a low-rank core space with the consistent topological structure. This framework enables FLoRA to effectively preserve the structural integrity of the original parameter space, such as maintaining the spatial locality in convolutional operations.
To tackle this challenge, recent works Chen et al. (2024a); Guo et al. (2020); He et al. (2021a); Hu et al. (2021a) have focused on the concept of parameter-efficient fine-tuning (PEFT), aiming to minimize the number of adjustable parameters required while achieving optimal task performance. These works specifically explore methods to model the incremental update of pre-trained weights in a manner that is economical in terms of parameters, without necessitating changes to the model’s architecture Zaken et al. (2021); Guo et al. (2020); Hu et al. (2021a). Among these works, LoRA Hu et al. (2021a) emerges as a pioneering effort, which proposes to adopt an additional term with a low-rank structure to the original weight. Specifically, the original weight matrix $\mathbf{W}_{0}\in\mathbb{R}^{d_{1}\times d_{2}}$ remains frozen, while a learnable low-rank $\Delta\mathbf{W}$ is added to $\mathbf{W}_{0}$, with the form as
@@
where $\mathbf{A}\in\mathbb{R}^{d_{2}\times r}$ and $\mathbf{B}\in\mathbb{R}^{d_{1}\times r}$ with $r\ll\{d_{1},d_{2}\}$. Given that $r$ is significantly smaller than the dimension of $\mathbf{W}_{0}$, LoRA necessitates the updating of only a limited number of trainable parameters, while achieving comparable performances. Subsequent to LoRA, recent studies He et al. (2021a); Bershatsky et al. (2024); Zhang et al. (2022a) endeavor to explore more efficient low-rank matrix adaptation methods concerning the matrix $\Delta\mathbf{W}$.
However, we have observed an intriguing phenomenon: a significant portion of existing works is narrowly focused on two-dimensional parameter spaces (i.e., linear layers), neglecting the existence of other dimensional spaces such as 2D convolution (4-dimension) layers, etc. But in practice, not all layers are linear across various models for downstream tasks. For example, ConvNeXt Liu et al. (2022) and Stable Diffusion Carreira & Zisserman (2017) are two models predominantly utilizing 2D convolutional layers. Alternatively, some approaches can be adapted for high-dimensional parameter spaces by directly adopting low-rank matrix adaptation. They models alternations in high-dimensional spaces into two dimensions, while neglecting the structural complexities of the original spaces. For instance, as detailed in Sec. 2.1 and Fig. 1, LoRA Hu et al. (2021a) compresses the changes in convolutions, which is a four-dimensional parameter space, into two dimensions. It subsequently applies low-rank matrix adaptation to the two-dimension space, intending to represent the alterations of the original four-dimensional parameter space. Yet, as discussed in Sec. 2.2, this adaptation fails to capture the inherent complexity and spatial locality specific to convolution operations. The result is a reshaped two-dimensional structure that compromises the integrity of the original parameter space, leading to a representation that does not fully encapsulate the change of convolutional space.
To this end, in this paper, we propose a low-rank tensor adaptation based method, FLoRA, represented as Fundamental LOw-Rank Adaptation. Positioned as a superior alternative to LoRA, FLoRA meets the ensuing three properties:
  • It can identify an appropriate low-rank representation for the changes in various dimensional parameter spaces, without destructing the structural integrity of the original parameter spaces.
@@
  • When applied to linear weights, with the same parameter budget, it requires similar training time and resources to that of LoRA, yet yield superior performance.
@@ 2.1 Low Rank Adaptation
LoRA Hu et al. (2021a) models the incremental update of a pre-trained weight matrix $\mathbf{W}_{0}\in\mathbb{R}^{d_{1}\times d_{2}}$ by the product of two small matrices $\mathbf{A}\in\mathbb{R}^{d_{2}\times r}$ and $\mathbf{B}\in\mathbb{R}^{d_{1}\times r}$, where $r\ll\{d_{1},d_{2}\}$. For $\mathbf{h}=\mathbf{W}_{0}\mathbf{x}$, the modified forward pass is
@@
where $\mathbf{A}_{c}\in\mathbb{R}^{kd_{out}\times r}$ and $\mathbf{B}_{c}\in\mathbb{R}^{kd_{in}\times r}$. Here, $Reshape(\mathbf{B}_{c}\mathbf{A}_{c}^{\mathsf{T}},\mathcal{W}_{c})$ involves altering the dimensions of $\mathbf{B}_{c}\mathbf{A}_{c}^{\mathsf{T}}$ to match those of $\mathcal{W}_{c}$. It is obvious that LoRA unfolds the original 4-dimensional parameter space $\Delta\mathcal{W}$ into a 2-dimensional space, subsequently leveraging the low-rank approximation of this 2-dimensional space to represent the original 4-dimensional construct. As shown later, this low-rank matrix adaptation method will destruct the structural integrity of the convolution layer.
2.2 Why Matrix Adaptation Breaks the Structural Integrity of the Convolution?
The changes of a high-dimension tensor such as convolutional parameter space can be modeled based on low-rank matrix adaptation following Eq. (3). However, during the reshaping process, elements that are adjacent within the convolutional kernel are those dispersed across various positions in the original matrix. More specifically, elements that are now localized within the convolutional kernel were located across multiple rows or columns of the matrix for reshaping. This shift poses significant challenges in learning spatial correlations among elements positioned disparately. Therefore, such a transformation disrupts the principle of locality inherent in the original convolutional operation, where each output element is determined by a small region of the input.
@@ 3 Method
In this section, we first introduce the formulation of FLoRA in an N-dimensional parameter space. Specifically, for a pre-trained weight $\mathcal{W}_{0}\in\mathbb{R}^{I_{1}\times I_{2}\times\cdots\times I_{N}}$ with N-dimension, we update the weights $\mathcal{W}_{0}$ with the changes $\Delta\mathcal{W}$ as
@@
without loss of generality. Here $\mathcal{G}\in\mathbb{R}^{J_{1}\times J_{2}\times\cdots\times J_{N}}$ and $\mathbf{A}^{(n)}\in\mathbb{R}^{I_{n}\times J_{n}}$, where $J_{n}\ll I_{n}$. $J_{1},...,J_{N}$ are hyper-parameters, and ×$\times_{n}$ denotes the mode-$n$ product between a tensor and a matrix. FLoRA considers the tensor $\mathcal{G}$ as a low-rank core space with the consistent topological structure to the original parameter space, with $\mathbf{A}^{(n)}$ representing the weights associated with each dimension. We then scale $\Delta\mathcal{W}$ by $s$ with $s$ being a constant. In the subsequent subsections, we will detail its manifestations in convolution and linear layers.
@@ 3.1 FLoRA for Convolution Layer
where $\mathcal{G}\in\mathbb{R}^{r_{1}\times r_{2}\times r_{3}\times r_{3}}$, $\mathbf{A}\in\mathbb{R}^{d_{in}\times r_{1}}$, $\mathbf{B}\in\mathbb{R}^{d_{out}\times r_{2}}$ and $\mathbf{C},\mathbf{D}\in\mathbb{R}^{k\times r_{3}}$. $r_{1}$ and $r_{2}$ are two hyper-parameters which are significantly smaller than $\{d_{in},d_{out}\}$. $r_{3}$ is another hyper-parameter that smaller than the kernel size $k$ of a convolution layer. Given that $3\times 3$ convolution is a prevalent configuration in convolutional foundation models Woo et al. (2023); Rombach et al. (2022); Wang et al. (2023a), $r_{3}$ is consequently chosen from {1,2}.
The core tensor $\mathcal{G}$ in FLoRA can be viewed as a compressed convolution parameter space. In essence, it serves as a core space for convolution. This means that in any convolution layer, there exists a convolution core, and what FLoRA aims to do is to determine this convolution core for each convolution space and configure the corresponding weights $\mathbf{A}$, $\mathbf{B}$, $\mathbf{C}$ and $\mathbf{D}$ for reconstructing that space. Different to low-rank matrix adaptation based methods, FLoRA does not need to alter the structure of the convolution. Alternatively, by learning the convolution core, FLoRA effectively preserves the convolution’s property of spatial locality.
Furthermore, while preserving or potentially augmenting the representational power of the convolution process, FLoRA achieves a remarkable reduction in the number of trainable parameters in comparison to LoRA. Assuming that the rank for both the input and output dimensions is uniform ($r_{1}=r_{2}=r$), the parameter requirement for FLoRA is calculated as $O(r_{3}^{2}r^{2}+r(d_{in}+d_{out})+2r_{3}k)$ parameters, while LoRA necessitate training at least $O(rk(d_{in}+d_{out}))$ parameters. Given that typically $r,k\ll\{d_{in},d_{out}\}$, therefore, FLoRA exhibits better parameter efficacy than LoRA as the number of the kernel increases.
@@ 4.1 Models and Datasets
Specifically, for CV tasks, we employ FLoRA to fine-tune the ConvNeXt-V2-L Woo et al. (2023), evaluating it on MS COCO Lin et al. (2014) by using Mask R-CNN He et al. (2017) implemented in MMDetection Chen et al. (2019), and on remote sensing image datasets DOTA Xia et al. (2018) with Oriented R-CNN Xie et al. (2021) based on MMRotate Zhou et al. (2022), and on the ADE20K Zhou et al. (2017) dataset thanks to UperNet Xiao et al. (2018) integrated in MMSegmentation Contributors (2020). We also employ FLoRA to fine-tune large-scale vision foundation model, i.e. InternViT-6B Chen et al. (2023), on ADE20K datasets. Detailed hyper-parameter settings can be find in Table 9 in Appendix.
For NLP tasks, we evaluate the DeBERTaV3-base He et al. (2021b) with FLoRA on the General Language Understanding Evaluation (GLUE) Wang et al. (2018) benchmark, which includes two single-sentence classification, three similarity and paraphrase and four natural language inference datasets. More details on GLUE dataset can be found in Table 10 in the Appendix.
For multi-modal tasks, we employ FLoRA to fine-tune LLaVA-1.5-7B Liu et al. (2024a), which is composed of a language model, Vicuna-1.5-7B Peng et al. (2023) and a vision encoder, CLIP ViT-L/336px Radford et al. (2021), on visual instruction tuning tasks, which include seven vision-language benchmarks: VQAv2 Goyal et al. (2017), GQA Hudson & Manning (2019), VisWiz Gurari et al. (2018), SQA Lu et al. (2022), VQAT Singh et al. (2019), POPE Li et al. (2023), and MMBench Liu et al. (2023).
Moreover, we fine-tune SDXLRombach et al. (2022) with FLoRA for generation task in Appendix A.1, LLaMA3-8B AI@Meta (2024) for commonsense reasoning tasks in Appendix A.2, and ViT-B Dosovitskiy (2020) for VTAB-1K benchmark Zhai et al. (2019) in Appendix A.3.